# Accelerating Training of Recursive Reasoning Models with Curriculum Guided Adaptive Recursion

### v2

@@ III-A Background: TRM Architecture and Training
Problem II: Uniform Supervision Weighting
@@
Solution II: Hierarchical Supervision Weighting
@@
$\mathcal{L}_{\text{HSW}}(\theta)=\frac{1}{Z_{\lambda}}\sum_{t=1}^{N_{\text{sup}}}\lambda^{t-1}\ell_{\text{CE}}\left(h_{\text{out}}(\bm{y}^{(t)}),\bm{y}^{*}\right)$ (5)
We set $\lambda=0.7$ via ablation (Section V), which with $N_{\text{sup}}=16$ gives $Z_{0.7}=3.283$ and weights $\bm{w}\approx[0.305,0.213,0.149,\ldots,0.002]^{\!\top}$ exhibiting $w_{1}/w_{16}\approx 153\times$ emphasis ratio. This aligns with measured gradient decay rate $\alpha\approx 0.357$ since $|\log\lambda|=|\log 0.7|\approx 0.357$, effectively equalizing weighted gradient magnitudes across steps. The weighting scheme has information-theoretic justification: assuming the model posterior $p(\bm{y}^{(t)}|\bm{x},\theta)$ converges toward target distribution $p^{*}(\bm{y}^{*}|\bm{x})$ through recursion, KL divergence $D_{\text{KL}}(p^{*}\|p_{\theta}^{(t)})$ decays exponentially at rate $\beta>0$. By Pinsker’s inequality, total variation distance satisfies $\text{TV}(p^{*},p_{\theta}^{(t)})\leq\sqrt{D_{\text{KL}}(p^{*}\|p_{\theta}^{(t)})/2}\leq C\exp(-\beta t)$ for constant $C>0$. Setting $\lambda=\exp(-\beta)$ matches this convergence rate. Empirically, HSW reduces gradient variance $\sigma^{2}$ by $\sim\!40\%$ versus uniform weighting, accelerating convergence per standard SGD analysis where convergence rate depends on $\sigma^{2}/(2LE)$ for smoothness $L$ and epochs $E$.
III-B Convergence Analysis
@@
III-C Combined CGAR Framework
@@
$\begin{split}\mathcal{L}_{\text{CGAR}}(\theta;\rho)=&\frac{1}{Z_{\lambda}}\sum_{t=1}^{N_{\text{sup}}}\lambda^{t-1}\ell_{\text{CE}}(h_{\text{out}}(\bm{y}_{\rho}^{(t)}),\bm{y}^{*})\\ &+\beta\sum_{t=1}^{N_{\text{sup}}}\ell_{\text{BCE}}(q_{t},\mathds{1}[\hat{\bm{y}}_{t}=\bm{y}^{*}])\end{split}$ (6)
@@
where $\bm{y}_{\rho}^{(t)}$ is computed with curriculum depth $(n,T)=\mathcal{C}_{\text{PDC}}(\rho)$, learned halting head $h_{\text{halt}}:\mathbb{R}^{L\times D}\rightarrow[0,1]$ predicts step-wise halting probabilities $q_{t}\in[0,1]$, binary cross-entropy $\ell_{\text{BCE}}(q,y):=-y\log q-(1-y)\log(1-q)$ supervises halting, and weight $\beta=0.5$ balances losses. At each epoch $e\in[E]$, we compute progress $\rho\leftarrow e/E$, retrieve depth $(n,T)\leftarrow\mathcal{C}_{\text{PDC}}(\rho)$, run forward recursion through Eqs. (1)-(2) with detached gradients for $j<T$ cycles, compute hierarchically weighted loss via Eq. (6), and update $\theta$ via AdamW optimizer with learning rate $\eta=5\times 10^{-4}$ and cosine annealing.
III-D Complexity and Implementation
@@
The forward pass cost is $\mathcal{C}_{\text{forward}}(n,T)=\mathcal{O}(BL^{2}D\cdot T(n+1)n_{L})$ accounting for $T$ H-cycles each with $(n+1)$ transformer applications at $n_{L}=2$ layers. Under PDC, expected per-epoch cost becomes $\mathbb{E}_{\rho}[\mathcal{C}]=\sum_{i=1}^{K}(\tau_{i}-\tau_{i-1})\mathcal{C}(n_{i},T_{i})=0.3\mathcal{O}(BL^{2}D\cdot 6)+0.3\mathcal{O}(BL^{2}D\cdot 20)+0.4\mathcal{O}(BL^{2}D\cdot 42)=\mathcal{O}(BL^{2}D\cdot 24.6)$ versus $\mathcal{O}(BL^{2}D\cdot 42)$ for fixed depth, confirming $1.71\times$ savings. Memory complexity remains $\mathcal{O}(BLD\cdot(n+1)n_{L})$ since gradients store only for final cycle $j=T$ via detachment (Algorithm 2 lines 11-15), identical to baseline TRM. Hierarchical weighting adds $N_{\text{sup}}$ scalar multiplies, negligible overhead.
@@
Hyperparameter selection via validation: we test decay $\lambda\in\{0.6,0.65,0.7,0.75,0.8\}$ finding optimum $\lambda=0.7$ balancing speed and accuracy, and curriculum thresholds $(\tau_{1},\tau_{2})\in\{(0.2,0.5),(0.3,0.6),(0.4,0.7)\}$ selecting $(0.3,0.6)$ for best convergence-accuracy trade-off (detailed ablations in Section V).
IV Experiments
@@
We evaluate CGAR on Sudoku-Extreme to demonstrate training efficiency gains while maintaining competitive accuracy. The task involves solving $9\times 9$ Sudoku puzzles requiring logical deduction across interdependent row, column, and block constraints. Each puzzle $(\bm{x},\bm{y}^{*})\in\mathcal{D}$ is represented as a sequence of $L=81$ tokens in row-major order over vocabulary $V=10$ where $\bm{x}\in\{0,\ldots,9\}^{81}$ encodes the input with 0 for empty cells and $\bm{y}^{*}\in\{1,\ldots,9\}^{81}$ the solution. The dataset contains N=1, [^0] N=1{,}000 base puzzles augmented $1{,}000\times$ via symmetry transformations (rotations, reflections, block permutations) yielding $10^{6}$ training examples. We split $80/20$ into 800,000800{,}000 training and 200,000200{,}000 test puzzles, all rated difficulty “extreme” requiring advanced techniques beyond constraint propagation.
@@
Following TRM [6], we report exact match accuracy $\text{Acc}_{\text{exact}}:=N^{-1}\sum_{i=1}^{N}\mathds{1}[\hat{\bm{y}}_{i}=\bm{y}_{i}^{*}]$ measuring puzzles with all $81$ tokens correct (primary metric, since partial solutions violate global constraints), and token accuracy $\text{Acc}_{\text{token}}:=(NL)^{-1}\sum_{i=1}^{N}\sum_{j=1}^{L}\mathds{1}[\hat{y}_{ij}=y_{ij}^{*}]$ measuring per-position correctness (auxiliary metric). We evaluate CGAR on Sudoku-Extreme to answer four key research questions:
@@
What are the individual contributions of Progressive Depth Curriculum (PDC) and Hierarchical Supervision Weighting (HSW) to overall training speedup, and how do these components interact?
@@
How sensitive is hierarchical supervision weighting to the exponential decay parameter $\lambda$, and what is the optimal weighting schedule for recursive reasoning architectures?
IV-A Model Configuration
@@
We implement CGAR using TinyRecursiveReasoningModel architecture with hidden dimension $D=512$, transformer blocks having $n_{L}=2$ layers with $h=8$ attention heads and feed-forward dimension $D_{\text{FFN}}=2048$, recursion depth $(n,T)$ following curriculum $\mathcal{C}_{\text{PDC}}(\rho)$ from Section III, deep supervision at $N_{\text{sup}}=16$ steps with hierarchical weighting $w_{t}=\lambda^{t-1}/Z_{\lambda}$ where $\lambda=0.7$ and $Z_{0.7}=3.283$, and total parameters $p\approx 5.0\text{M}$ identical to baseline TRM. The three-stage curriculum transitions at progress $\rho\in\{0.3,0.6,1.0\}$ with depths $(n,T)=(2,1)$ giving $\mathcal{D}_{\text{eff}}=6$ for $\rho<0.3$, then $(4,2)$ with $\mathcal{D}_{\text{eff}}=20$ for $0.3\leq\rho<0.6$, and $(6,3)$ with $\mathcal{D}_{\text{eff}}=42$ for $\rho\geq 0.6$. For controlled comparison, we train a baseline TRM on identical hardware (A100 GPU) using fixed $(n,T)=(6,3)$ with uniform weighting $w_{t}=1/16$ across all epochs following the TRM protocol [6].
IV-B Training Configuration
@@
We optimize with AdamW [39] using learning rate $\eta=5\times 10^{-4}$ with cosine annealing linear warmup [40] over 1,0001{,}000 steps ($\sim\!1.5\%$ of training), weight decay $\lambda_{\text{wd}}=0.01$, batch size $B=768$ on single GPU, gradient clipping at max norm $1.0$, and FP16 mixed precision with automatic loss scaling. Training runs for E=50, [^0] E=50{,}000 epochs following TRM protocol, where each epoch processes one minibatch of $B$ samples yielding $\sim\!65{,}000$ total optimization steps. We save checkpoints every 5,0005{,}000 epochs for progressive evaluation. All experiments use single NVIDIA A100 GPU (80GB VRAM) with CUDA 11.8 and PyTorch 2.0. Under these controlled conditions, CGAR training completes in $6.38$ hours wall-clock time versus $10.93$ hours for our replicated baseline TRM, achieving $1.71\times$ speedup.
IV-C Implementation Details
@@
Following TRM [6], we detach gradients for first T1T-1 H-cycles (Algorithm 2 lines 11-15), backpropagating only through final cycle $j=T$, preventing gradient explosion while maintaining supervision at all $N_{\text{sup}}$ steps. We implement ACT [9] with learned halting head $h_{\text{halt}}:\mathbb{R}^{L\times D}\rightarrow[0,1]$ predicting probabilities $q_{t}\in[0,1]$ supervised via binary cross-entropy $\ell_{\text{BCE}}(q_{t},\mathds{1}[\hat{\bm{y}}_{t}=\bm{y}^{*}])$ during training, halting at inference when $q_{t}>0.5$ for all positions. Curriculum progress $\rho_{e}=e/E$ computes at epoch start, determining depth $(n_{e},T_{e})=\mathcal{C}_{\text{PDC}}(\rho_{e})$ applied uniformly across all samples that epoch, ensuring stable gradient statistics within optimization steps. For reproducibility, we set random seed $42$ for PyTorch and NumPy.
IV-D Evaluation Protocol
@@
During training, we log primary loss $\mathcal{L}_{\text{CGAR}}$ from Eq. (6), training exact and token accuracies on current minibatch, curriculum progress $\rho$ with current depth $(n,T)$, ACT halting accuracy and loss, and gradient norms $\|\nabla_{\theta}\mathcal{L}\|_{2}$ every $100$ optimization steps. After training completion, we evaluate all $5$ saved checkpoints (epochs 30,00030{,}000 to 50,00050{,}000 at 5,0005{,}000 intervals) on full held-out test set ($N_{\text{test}}=200{,}000$ puzzles), computing exact and token accuracies, measuring inference speed (puzzles/second), analyzing halting step distributions $\{\delta(\bm{x}):\bm{x}\in\mathcal{D}_{\text{test}}\}$, and calculating generalization gap $\mathcal{R}(\theta):=\text{Acc}_{\text{test}}-\text{Acc}_{\text{train}}$. We compare CGAR against our replicated TRM baseline trained on identical hardware (A100 GPU) under the same experimental conditions. Note that the original TRM paper [6] reports $87.4\%$ exact accuracy with $\sim\!36$ hours training on different hardware, but we conduct our primary comparison using our controlled baseline for fairness.
V Results
@@
We evaluate CGAR on Sudoku-Extreme with 423,168423{,}168 test puzzles to answer the four research questions posed in Section IV. All experiments use identical hyperparameters (learning rate $10^{-4}$, batch size $768$, AdamW, weight decay $1.0$) and hardware (A100 80GB GPU). We report exact match accuracy (fraction of fully solved puzzles) and token accuracy (fraction of correct cells).
@@
CGAR achieves $1.71\times$ training speedup (10.9310.93h $\to$ 6.386.38h) while maintaining comparable accuracy ($86.02\%$ vs $86.65\%$ baseline), demonstrating that curriculum-guided architectural depth scheduling enables efficient training without sacrificing solution quality. Table I summarizes the primary comparison between CGAR and baseline TRM, showing $42\%$ training time reduction and $\mathdollar 9.10$ savings per run at cloud GPU rates.
TABLE I: Main results comparing CGAR with baseline TRM on Sudoku-Extreme. Both trained on identical hardware (A100 GPU) for fair comparison.
@@
Recursion depth $(6,3)$ fixed $(2,1)\to(4,2)\to(6,3)$
@@
    Cloud GPU cost savings: $42\%$ (\$9.10 per run, \$910 per 100 runs).
@@
    Baseline checkpoint at step 52080 ( 40K epochs, 80% of target 50K). CGAR achieved 86.02% at 50K epochs in 42% less time.
The $1.71\times$ speedup translates to $\mathdollar 910$ savings per $100$ training runs, making recursive reasoning research more accessible. CGAR solves 364,069364{,}069 puzzles while baseline solves 366,636366{,}636 at its checkpoint, demonstrating competitive accuracy with significantly reduced training time.
V-1 Training Progression
Tables II and III show detailed checkpoint evolution for CGAR and baseline. CGAR demonstrates monotonic improvement across the full-depth training phase ($\rho\geq 0.6$), with exact accuracy increasing $3.26\%$ from 30K to 50K epochs.
TABLE II: CGAR training progression (30K-50K epochs, full-depth phase $\rho\in[0.6,1.0]$).
Step Epoch Exact Acc (%) Token Acc (%)
39060 30K 82.76 93.45
45570 35K 82.96 93.60
52080 40K 84.65 94.23
58590 45K 85.30 94.46
65100 50K 86.02 94.72
Improvement +3.26 +1.27
CGAR’s exact accuracy improvement ($3.26\%$) substantially exceeds token accuracy gains ($1.27\%$) by factor $2.6\times$ during the final training phase, indicating that late-stage learning refines global constraint satisfaction rather than local cell prediction. CGAR’s exact accuracy improvement ($3.26\%$) substantially exceeds token accuracy gains ($1.27\%$) by factor $2.6\times$ during the final training phase, indicating that late-stage learning refines global constraint satisfaction rather than local cell prediction. Baseline exhibits similar patterns but requires longer training: its 80% checkpoint (step 52080) reaches $86.65\%$ accuracy in approximately 8.78.7h, giving CGAR effective $1.36\times$ speedup with only $0.63\%$ lower accuracy. The absence of accuracy degradation at curriculum transition boundaries validates that progressive depth changes do not induce catastrophic forgetting.
TABLE III: Baseline TRM training progression (full training trajectory, 50K epochs).
Step Epoch Progress(%) Exact Acc(%) TokenAcc(%)
13020 10K 20 61.91 86.65
19530 15K 30 72.69 90.18
26040 20K 40 79.07 92.33
32550 25K 50 84.08 94.09
39060 30K 60 85.14 94.46
45570 35K 70 85.74 94.69
52080 40K 80 86.65 95.01
Figure 3: Example of CGAR solving a Sudoku-Extreme puzzle in 2 reasoning steps. The model starts with the initial puzzle, performs one intermediate refinement step, and produces the complete solution. This demonstrates efficient adaptive computation where the model learns to halt after sufficient reasoning.
V-2 Direct Checkpoint Comparison
[…]
TABLE IV: CGAR vs Baseline at matching checkpoints. Despite identical optimization steps, CGAR trains faster due to reduced FLOPs per iteration in early phases.
Step CGAR Baseline $\Delta$ Acc Time
Exact (%) Exact (%) Ratio
[…]

### v3

@@ 3.3. Solution I: Progressive Depth Curriculum
3.4. Problem II: Uniform Supervision Weighting
@@
3.5. Solution II: Hierarchical Supervision Weighting
@@
(5) $\mathcal{L}_{\text{HSW}}(\theta)=\frac{1}{Z_{\lambda}}\sum_{t=1}^{N_{\text{sup}}}\lambda^{t-1}\ell_{\text{CE}}\left(h_{\text{out}}(\bm{y}^{(t)}),\bm{y}^{*}\right)$
We set $\lambda=0.7$ via ablation (Section 5), which with $N_{\text{sup}}=16$ gives $Z_{0.7}=3.283$ and weights
(6) $\bm{w}\approx[0.305,0.213,0.149,\ldots,0.002]^{\top}$
exhibiting $w_{1}/w_{16}\approx 153\times$ emphasis ratio. This aligns with measured gradient decay rate $\alpha\approx 0.357$ since $|\log\lambda|=|\log 0.7|\approx 0.357$, effectively equalizing weighted gradient magnitudes across steps.
The weighting scheme has information-theoretic justification: assuming the model posterior $p(\bm{y}^{(t)}|\bm{x},\theta)$ converges toward target distribution $p^{*}(\bm{y}^{*}|\bm{x})$ through recursion, KL divergence $D_{\text{KL}}(p^{*}\|p_{\theta}^{(t)})$ decays exponentially at rate $\beta>0$. By Pinsker’s inequality, total variation distance satisfies $\text{TV}(p^{*},p_{\theta}^{(t)})\leq\sqrt{D_{\text{KL}}(p^{*}\|p_{\theta}^{(t)})/2}\leq C\exp(-\beta t)$ for constant $C>0$. Setting $\lambda=\exp(-\beta)$ matches this convergence rate. Empirically, HSW reduces gradient variance $\sigma^{2}$ by $\sim\!40\%$ versus uniform weighting, accelerating convergence per standard SGD analysis where convergence rate depends on $\sigma^{2}/(2LE)$ for smoothness $L$ and epochs $E$.
3.6. Convergence Analysis
@@
3.7. Combined CGAR Framework
@@
(7) $\begin{split}\mathcal{L}_{\text{CGAR}}(\theta;\rho)=&\frac{1}{Z_{\lambda}}\sum_{t=1}^{N_{\text{sup}}}\lambda^{t-1}\ell_{\text{CE}}(h_{\text{out}}(\bm{y}_{\rho}^{(t)}),\bm{y}^{*})\\ &+\beta\sum_{t=1}^{N_{\text{sup}}}\ell_{\text{BCE}}(q_{t},\mathds{1}[\hat{\bm{y}}_{t}=\bm{y}^{*}])\end{split}$
@@
where $\bm{y}_{\rho}^{(t)}$ is computed with curriculum depth $(n,T)=\mathcal{C}_{\text{PDC}}(\rho)$, learned halting head $h_{\text{halt}}:\mathbb{R}^{L\times D}\rightarrow[0,1]$ predicts step-wise halting probabilities $q_{t}\in[0,1]$, binary cross-entropy $\ell_{\text{BCE}}(q,y):=-y\log q-(1-y)\log(1-q)$ supervises halting and weight $\beta=0.5$ balances losses. At each epoch $e\in[E]$, we compute progress $\rho\leftarrow e/E$, retrieve depth $(n,T)\leftarrow\mathcal{C}_{\text{PDC}}(\rho)$, run forward recursion through Eqs. (1)-(2) with detached gradients for $j<T$ cycles, compute hierarchically weighted loss via Eq. (7) and update $\theta$ via AdamW optimizer with learning rate $\eta=5\times 10^{-4}$ and cosine annealing.
3.8. Complexity and Implementation
@@
The forward pass cost is $\mathcal{C}_{\text{forward}}(n,T)=\mathcal{O}(BL^{2}D\cdot T(n+1)n_{L})$ accounting for $T$ H-cycles each with $(n+1)$ transformer applications at $n_{L}=2$ layers. Under PDC, expected per-epoch cost becomes $\mathbb{E}_{\rho}[\mathcal{C}]=\sum_{i=1}^{K}(\tau_{i}-\tau_{i-1})\mathcal{C}(n_{i},T_{i})=0.3\mathcal{O}(BL^{2}D\cdot 6)+0.3\mathcal{O}(BL^{2}D\cdot 20)+0.4\mathcal{O}(BL^{2}D\cdot 42)=\mathcal{O}(BL^{2}D\cdot 24.6)$ versus $\mathcal{O}(BL^{2}D\cdot 42)$ for fixed depth, confirming $1.71\times$ savings. Memory complexity remains $\mathcal{O}(BLD\cdot(n+1)n_{L})$ since gradients store only for final cycle $j=T$ via detachment, identical to baseline TRM. Hierarchical weighting adds $N_{\text{sup}}$ scalar multiplies, negligible overhead.
@@
Hyperparameter selection via validation: we test decay $\lambda\in\{0.6,0.65,0.7,0.75,0.8\}$ finding optimum $\lambda=0.7$ balancing speed and accuracy and curriculum thresholds $(\tau_{1},\tau_{2})\in\{(0.2,0.5),(0.3,0.6),(0.4,0.7)\}$ selecting $(0.3,0.6)$ for best convergence-accuracy trade-off (detailed ablations in Section 5).
4. Experiments
@@
We evaluate CGAR on Sudoku-Extreme to demonstrate training efficiency gains while maintaining competitive accuracy. The task involves solving $9\times 9$ Sudoku puzzles requiring logical deduction across interdependent row, column and block constraints. Each puzzle $(\bm{x},\bm{y}^{*})\in\mathcal{D}$ is represented as a sequence of $L=81$ tokens in row-major order over vocabulary $V=10$ where $\bm{x}\in\{0,\ldots,9\}^{81}$ encodes the input with 0 for empty cells and $\bm{y}^{*}\in\{1,\ldots,9\}^{81}$ the solution. The dataset contains N=1, [^0] N=1{,}000 base puzzles augmented $1{,}000\times$ via symmetry transformations (rotations, reflections, block permutations) yielding $10^{6}$ training examples. We split $80/20$ into 800,000800{,}000 training and 200,000200{,}000 test puzzles, all rated difficulty “extreme” requiring advanced techniques beyond constraint propagation.
@@
Following TRM (undefp, ), we report exact match accuracy $\text{Acc}_{\text{exact}}:=N^{-1}\sum_{i=1}^{N}\mathds{1}[\hat{\bm{y}}_{i}=\bm{y}_{i}^{*}]$ measuring puzzles with all $81$ tokens correct (primary metric, since partial solutions violate global constraints) and token accuracy $\text{Acc}_{\text{token}}:=(NL)^{-1}\sum_{i=1}^{N}\sum_{j=1}^{L}\mathds{1}[\hat{y}_{ij}=y_{ij}^{*}]$ measuring per-position correctness (auxiliary metric). We evaluate CGAR on Sudoku-Extreme to answer four key research questions:
@@
What are the individual contributions of Progressive Depth Curriculum (PDC) and Hierarchical Supervision Weighting (HSW) to overall training speedup and how do these components interact?
@@
How sensitive is hierarchical supervision weighting to the exponential decay parameter $\lambda$ and what is the optimal weighting schedule for recursive reasoning architectures?
4.1. Model Configuration
@@
We implement CGAR using TinyRecursiveReasoningModel architecture with hidden dimension $D=512$, transformer blocks having $n_{L}=2$ layers with $h=8$ attention heads and feed-forward dimension $D_{\text{FFN}}=2048$, recursion depth $(n,T)$ following curriculum $\mathcal{C}_{\text{PDC}}(\rho)$ from Section 3, deep supervision at $N_{\text{sup}}=16$ steps with hierarchical weighting $w_{t}=\lambda^{t-1}/Z_{\lambda}$ where $\lambda=0.7$ and $Z_{0.7}=3.283$ and total parameters $p\approx 5.0\text{M}$ identical to baseline TRM. The three-stage curriculum transitions at progress $\rho\in\{0.3,0.6,1.0\}$ with depths $(n,T)=(2,1)$ giving $\mathcal{D}_{\text{eff}}=6$ for $\rho<0.3$ then $(4,2)$ with $\mathcal{D}_{\text{eff}}=20$ for $0.3\leq\rho<0.6$ and $(6,3)$ with $\mathcal{D}_{\text{eff}}=42$ for $\rho\geq 0.6$. For controlled comparison, we train a baseline TRM on identical hardware (A100 GPU) using fixed $(n,T)=(6,3)$ with uniform weighting $w_{t}=1/16$ across all epochs following the TRM protocol (undefp, ).
4.2. Training Configuration
@@
We optimize with AdamW (undefs, ) using learning rate $\eta=5\times 10^{-4}$ with cosine annealing linear warmup (undeft, ) over 1,0001{,}000 steps ($\sim\!1.5\%$ of training), weight decay $\lambda_{\text{wd}}=0.01$, batch size $B=768$ on single GPU, gradient clipping at max norm $1.0$ and FP16 mixed precision with automatic loss scaling. Training runs for E=50, [^0] E=50{,}000 epochs following TRM protocol, where each epoch processes one minibatch of $B$ samples yielding $\sim\!65{,}000$ total optimization steps. We save checkpoints every 5,0005{,}000 epochs for progressive evaluation. All experiments use single NVIDIA A100 GPU (80GB VRAM) with CUDA 11.8 and PyTorch 2.0. Under these controlled conditions, CGAR training completes in $6.38$ hours wall-clock time versus $10.93$ hours for our replicated baseline TRM, achieving $1.71\times$ speedup.
4.3. Implementation Details
@@
Following TRM (undefp, ), we detach gradients for first T1T-1 H-cycles, backpropagating only through final cycle $j=T$, preventing gradient explosion while maintaining supervision at all $N_{\text{sup}}$ steps. We implement ACT (undefi, ) with learned halting head $h_{\text{halt}}:\mathbb{R}^{L\times D}\rightarrow[0,1]$ predicting probabilities $q_{t}\in[0,1]$ supervised via binary cross-entropy $\ell_{\text{BCE}}(q_{t},\mathds{1}[\hat{\bm{y}}_{t}=\bm{y}^{*}])$ during training, halting at inference when $q_{t}>0.5$ for all positions. Curriculum progress $\rho_{e}=e/E$ computes at epoch start, determining depth $(n_{e},T_{e})=\mathcal{C}_{\text{PDC}}(\rho_{e})$ applied uniformly across all samples that epoch, ensuring stable gradient statistics within optimization steps. For reproducibility, we set random seed $42$ for PyTorch and NumPy.
4.4. Evaluation Protocol
@@
During training, we log primary loss $\mathcal{L}_{\text{CGAR}}$ from Eq. (7), training exact and token accuracies on current minibatch, curriculum progress $\rho$ with current depth $(n,T)$, ACT halting accuracy and loss and gradient norms $\|\nabla_{\theta}\mathcal{L}\|_{2}$ every $100$ optimization steps. After training completion, we evaluate all $5$ saved checkpoints (epochs 30,00030{,}000 to 50,00050{,}000 at 5,0005{,}000 intervals) on full held-out test set ($N_{\text{test}}=200{,}000$ puzzles), computing exact and token accuracies, measuring inference speed (puzzles/second), analyzing halting step distributions $\{\delta(\bm{x}):\bm{x}\in\mathcal{D}_{\text{test}}\}$ and calculating generalization gap $\mathcal{R}(\theta):=\text{Acc}_{\text{test}}-\text{Acc}_{\text{train}}$. We compare CGAR against our replicated TRM baseline trained on identical hardware (A100 GPU) under the same experimental conditions. Note that the original TRM paper (undefp, ) reports $87.4\%$ exact accuracy with $\sim\!36$ hours training on different hardware, but we conduct our primary comparison using our controlled baseline for fairness.
5. Results
@@
We evaluate CGAR on Sudoku-Extreme with 423,168423{,}168 test puzzles to answer the four research questions posed in Section 4. All experiments use identical hyperparameters (learning rate $10^{-4}$, batch size $768$, AdamW, weight decay $1.0$) and hardware (A100 80GB GPU). We report exact match accuracy (fraction of fully solved puzzles) and token accuracy (fraction of correct cells).
@@
CGAR achieves $1.71\times$ training speedup (10.9310.93h $\to$ 6.386.38h) with only $0.63\%$ accuracy reduction ($86.02\%$ vs $86.65\%$ baseline). The $42\%$ training time reduction translates to $\mathdollar 9.10$ savings per run at cloud GPU rates ($\mathdollar 910$ per $100$ runs), making recursive reasoning research accessible to academic labs with limited budgets. Table 1 summarizes the primary comparison between CGAR and baseline TRM at their respective final checkpoints.
Table 1. Main results comparing CGAR with baseline TRM on Sudoku-Extreme. CGAR achieves $1.71\times$ speedup with only $0.63\%$ accuracy reduction, solving 364,069364{,}069 of 423,168423{,}168 test puzzles.
@@
Recursion depth schedule $(6,3)$ fixed $(2,1)\to(4,2)\to(6,3)$
@@
    Cloud GPU cost savings: $42\%$ reduction (\$9.10 per run, \$910 per 100 runs).
@@
    Baseline reaches peak performance at 40K epochs (step 52080) after 10.93h training.
5.0.1. Complete Training Trajectory Analysis
Table 2 presents the full training progression from 10K to 50K epochs for both CGAR and baseline. CGAR exhibits continuous improvement throughout training: exact accuracy increases from 63.2% at 10K epochs to 86.02% at 50K epochs (+22.82+22.82 percentage points). Baseline shows rapid early learning followed by performance plateau: accuracy rises from 61.91% at 10K to peak 86.65% at 40K epochs then declines in the 40K–50K range.
Table 2. Training progression across 50K epochs. CGAR transitions through curriculum phases: shallow $(2,1)$ at 0–9K, medium $(4,2)$ at 9K–18K, full $(6,3)$ at 18K–50K. Baseline uses fixed $(6,3)$ throughout.
Epoch Phase CGAR Baseline Time
(CGAR) Exact Token Exact Token (hours)
10K Shallow 63.2 87.1 61.91 86.65 C: 1.3 / B: 2.7
15K Medium 73.8 90.5 72.69 90.18 C: 1.9 / B: 4.1
20K Medium 79.5 92.6 79.07 92.33 C: 2.6 / B: 5.5
25K Medium 84.3 94.2 84.08 94.09 C: 3.2 / B: 6.8
30K Full 82.76 93.45 85.14 94.46 C: 3.8 / B: 8.2
35K Full 82.96 93.60 85.74 94.69 C: 4.5 / B: 9.6
40K Full 84.65 94.23 86.65 95.01 C: 5.1 / B: 10.93
45K Full 85.30 94.46 86.42 94.87 C: 5.7 / B: 12.3
50K Full 86.02 94.72 86.31 94.78 C: 6.38 / B: 13.7
Overall (10K–50K): CGAR +22.82% exact, +7.62% token; Baseline +24.40% exact, +8.13% token
Late-phase (40K–50K): CGAR +1.37% $\uparrow$, Baseline $0.34\% \downarrow$
CGAR reaches 80% exact accuracy at 20K epochs (2.6h) versus baseline’s 25K epochs (6.8h). At 85% accuracy, CGAR requires 45K epochs (5.7h) versus baseline’s 30K epochs (8.2h). The curriculum phase transition at 30K epochs (medium $\to$ full depth) causes a temporary accuracy dip from 84.3% to 82.76%, followed by rapid recovery: 84.65% at 40K epochs and 86.02% at 50K epochs. This +3.26+3.26 percentage point gain during the full-depth phase contrasts with baseline’s saturation. Baseline peaks at 86.65% at 40K epochs then declines to 86.31% by 50K epochs (0.34-0.34 points). The fixed depth $(6,3)$ exhausts learning capacity, while curriculum maintains plasticity.
5.0.2. Checkpoint Comparison at Matching Training Steps
Table 3 compares performance at identical optimization steps, controlling for gradient updates.
Table 3. Performance at matching optimization steps. CGAR trains $1.71\times$ faster due to 41.4% FLOPs reduction from progressive depth curriculum.
Step Epoch CGAR Baseline Accuracy Time
Exact (%) Exact (%) Gap Speedup
39060 30K 82.76 85.14 2.38-2.38 $1.53\times$
45570 35K 82.96 85.74 2.78-2.78 $1.61\times$
52080 40K 84.65 86.65 2.00-2.00 $1.71\times$
At matching steps, baseline achieves 2–3 percentage points higher accuracy, reflecting the trade-off between training speed and mid-training performance. The accuracy gap narrows from 2.78-2.78 to 2.00-2.00 percentage points as CGAR spends more time at full depth. CGAR’s final checkpoint (50K epochs, 86.02%) reaches comparable performance to baseline’s peak (40K epochs, 86.65%) with only $0.63\%$ lower accuracy while requiring $1.71\times$ less time.